How do I put this....(a+bi)^3.....in this answer form....
(a+bi)^3=______ + ______ i.

Answer by kapilsinghi(68)

(a+bi)^3
= a^3 + (bi)^3 + 3a(bi)^2 + 3bi(a)^2
= a^3 - ib^3 - 3a(b)^2 + i3b(a)^2
= a^3 - 3a(b)^2 + i3b(a)^2 - ib^3
= (a^3 - 3a(b)^2) + i(3b(a)^2 - b^3)
kapilsinghi123@gmail.com